1)     The average monthly salary paid to all employees in a company was Rs 8,000. The average monthly salaries paid to male and female employees of the company were Rs 10,600 and Rs 7,500 respectively. Find out the percentages of males and females employed by the company.


1
Expert's answer
2021-01-29T12:39:58-0500

Let "x=" the percentages of males and let "y=" the percentages of females. Then


"x+y=100"

"10600(\dfrac{x}{100})+7500(\dfrac{y}{100})=8000"

Solve


"106x+75(100-x)=8000"

"31x=500"

"x=\dfrac{500}{31}"

"y=100-\dfrac{500}{31}=\dfrac{2600}{31}"

"\dfrac{500}{31}\approx16.129"


"\dfrac{2600}{31}\approx83.871"

The percentages of males employed by the company is 16.129%.

The percentages of females employed by the company is 83.871%.